A recruiting firm finds that 20% of the applicants for a particularsales position are fluent in both English and Spanish. Applicantsare selected at random from a very large pool and interviewedsequentially. Suppose fours jobs requiring fluency in English andSpanish are open.Xis a random variable defined as the number of unqualifiedapplicants interviewed before finding the fourth qualified applicant.Yis a random variable defined as the total number of applicantsinterviewed before finding the fourth qualified applicant.What is the relationship betweenXandY?
Answer #1

since the probability is until the job is filled the distribution of this probability is geometric so

mean = 1 / p * cost

SD = srqt ( q / p^2 ) * cost

cost = $125

p = 0.20

q = 1-p = 1 - 0.20 = 0.80

expected value will be = 1/0.20 * 125 = 625

SD = srqt ( 0.80 / 0.20^2 ) * 125 = 559.017